OpenStudy (anonymous):

The slope of a line is -13/3. The point (-3, 7) is on the line. Determine the equation of this line

OpenStudy (anonymous):

13x+3y=-18

OpenStudy (katrinakaif):

Plug it intot he equation (y-y1) = m (x-x1) You arleady have M and you already have your coordinates that you can plug in as (x1,y1)

OpenStudy (anonymous):

\[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

13x+3y=-18

OpenStudy (anonymous):

x1=-3 and y1=7